OpenStudy (anonymous):

Solve: x2 + 2x - 1 = 0

Nnesha (nnesha):

i'm pretty sure you know how to factor that :-

Parth (parthkohli):

That can't be factored, honestly.

OpenStudy (anonymous):

i got \[-2\pm \sqrt{8}> divided by 2\]

OpenStudy (anonymous):

oh wait

Parth (parthkohli):

Looks alright to me. You can write that as \(-1 \pm \sqrt 2\)

Nnesha (nnesha):

factor 8

OpenStudy (anonymous):

square root of 8 is a long decimal

OpenStudy (anonymous):

\[x^2+2x=1,add~1 ~\to~both~sides\]

Parth (parthkohli):

\[\sqrt 8 = 2 \sqrt 2\]

Nnesha (nnesha):

\[\huge\rm \frac{ -2 \pm \sqrt{4 \times 2}}{ 2}\] yep right that's why you to factor 8 4 times 2 =8 now take square root of 4

Nnesha (nnesha):

have**

OpenStudy (anonymous):

thats what i got, but what next?

Parth (parthkohli):

Then you've solved it.

OpenStudy (anonymous):

1 + √2 and -1 - √2 -1 + √2 and -1 - √2 -1 + √3 and -1 + √2 -1 - √2 and 1 + √2

Nnesha (nnesha):

\[\huge\rm \frac{ -2 \pm 2\sqrt{ \times 2}}{ 2}\] 2 is common so take it out

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so,..

OpenStudy (anonymous):

i will fan all of u if you help me with 3 questions